What would be the R script to reshape a 6 ×24 numeric matrix, say tmp, into a 8 × 6 × 3 array?

A. matrix(tmp,8,6,3) B. array(matrix(tmp,8,6,3)) C. array(tmp, dim=c(8,6,3)) D. as.array(tmp,8,6,3) E. array(tmp, dim=c(3,6,8))
Answer #1

The answer to the above problem is as follows-

The correct option is C.array(tmp, dim=c(8,6,3))

The options along with explanation is as follows-

A.  matrix(tmp,8,6,3) - Wrong option

Reason - This statement changes the 8*24 matrix to a dimension of 8*6 matrix

B. array(matrix(tmp,8,6,3)) - Wrong option.

Reason - This statement changes the 8*24 matrix to an array of length 48.

C.array(tmp, dim=c(8,6,3)) - CORRECT ANSWER

Explanation - This statement changes the 8*24 matrix to a 8*6*3 dimensional array.

D. as.array(tmp,8,6,3) - Wrong option

Reason - This statement changes the 8*24 matrix to an array of dimension 8*24.So the dimension does not change.

E. array(tmp, dim=c(3,6,8)) - Wrong option.

Reason - This statement changes the 8*24 matrix to an array of dimension 3*6*8. And not 8*6*3. So this is also incorrect option.

If this answer helps, please give an up vote and feel free to comment for any query.